2log9x=log9(5x-4)








Answer by nerdybill(5404)

2log9x=log9(5x-4)
log9(x^2)=log9(5x-4)
x^2 = 5x-4
x^2-5x+4 = 0
(x-4)(x-1) = 0
.
x = {1,4}